OpenStudy (anonymous):

Probability Question! A game of chance costs $2.00 to play. If you draw a spade from a deck of cards you win $5.00. What is the expected value?

OpenStudy (anonymous):

I need help on solving it, not just getting the answer. So explain if possible?

OpenStudy (anonymous):

expected values is the amount you win times the probability you win it minus the amount you lose times the probability you lose it

OpenStudy (kropot72):

Let the dollar profit be X. For a win, the profit is $3. For a loss, the profit is -$2.  \[E(X)=(-2\times\frac{39}{52})+(3\times\frac{13}{52})=you\ can\ calculate\]